OpenStudy (anonymous):

Is the sequence 3, 12, 36, ... a geometric sequence? Explain.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

no, 12/3=4, but 36/12=3. since these quotients are not equal, there is not a common ratio.

OpenStudy (anonymous):

The definition of geom series is that the RATIO of adjacent numbers is CONSTANT

OpenStudy (anonymous):

so @madisonmaruhnich94 do you have this ratio constant here ?

OpenStudy (anonymous):

12/3 = ? and 36/12 = ?

OpenStudy (anonymous):

12/3=4 and 36/12=3

OpenStudy (anonymous):

so what doy conclude ?

OpenStudy (anonymous):

there not equal so there for there not a common ratio

OpenStudy (anonymous):

Yeap

OpenStudy (anonymous):

soo ?